OpenStudy (anonymous):

Evaluate the summation of 3 n plus 2, from n equals 1 to 14..

OpenStudy (anonymous):

\[\sum_{n=1}^{14} 3n + 2\]

OpenStudy (anonymous):

Answer options A) 39 B) 49 C) 340 D) 343

OpenStudy (johnweldon1993):

Well first we can break it up using properties of summation \[\sum_{i}^{n} (c + ai) = \sum_{i = 1}^{n}c + \sum_{i=1}^{n}ai\] So here, we would have \[\sum_{n=1}^{14} 3n + \sum_{n=1}^{14} 2\] Using the properties of summation we know that \[\sum_{i}^{n}c = cn\] and \[\sum_{i}^{n}i = \frac{n(n + 1)}{2}\] So lets rewrite yours as \[3\sum_{n=1}^{14} n + \sum_{n=1}^{14} 2\] Which would be \[\large 3\frac{14(15)}{2} + 2(14)\] What does that come out to?

OpenStudy (agreene):

\[\sum_{n=1}^{14}3n+2=\sum_{n=1}^{14}3n+\sum_{n=1}^{14}2=3\sum_{n=1}^{14}n+\sum_{n=1}^{14}2\] remember the rule \[\sum_{i=1}^{n}i=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

id I tried to answer it but a reaaaaaaaaally long number came up @johnweldon1993

OpenStudy (agreene):

do the algebra that john listed at the end and you should find a useful answer

OpenStudy (anonymous):

the number is long

OpenStudy (johnweldon1993):

We have \[\large \frac{3\times 14\times 15}{2} + 2(14)\]

OpenStudy (johnweldon1993):

So 3*14*15 = 630 That divided by 2 = 315 And 315 + 2(14) 2 times 14 = 28 315 + 28 = ?

OpenStudy (anonymous):

3433

OpenStudy (johnweldon1993):

...assuming you accidentally put an additional 3 at the end...

OpenStudy (anonymous):

yeah sorry lol